OpenStudy (anonymous):

Which quadratic function has an inverse? A) f(x) = x2 + 3, x ≤ 0 B) f(x) = x2 + 3, x ≤ 4 C) f(x) = x2 + 3, x ≥ -1 D) f(x) = x2 + 3, x ≥ ∞ And how to do it plz

OpenStudy (anonymous):

Answer is B

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

To find the inverse of a function you first have switch the domain and range

OpenStudy (anonymous):

y = x^2 + 3 switch the y and x to get x = y^2 + 3

OpenStudy (anonymous):

solve for x and you get \[y = \sqrt{x - 3}\] where \[x \ge 3\]

OpenStudy (anonymous):

so you solve and youll get the inverse

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

it cant be A or C because x≥3, and it cant be D because nothing is greater than infinity or equal to infinity other than infinity

OpenStudy (anonymous):

alright, thank you i did this a few months ago, probally why i dont remember, thanks again!